Shoe company A domestic shoe company distributes running shoes and tennis shoes...continues

Shoe company
A domestic shoe company distributes running shoes and tennis shoes for
$95 per pair. the marginal cost of producing a pair of running shoes
is $60, and the marginal cost of producing a pair of tennis shoes is
$45. A Chinese retailer offers to purchase running shoes for $55 per
pair and tennis shoes for $55 per pair for distribution in China.
Should the shoe company sell any shoes to the Chinese retailer?
(Ignore any potential issues of bundling the two type of shoes
together as part of the sale and any competitive effects that
international sales might have on current domestic sales.)
Answer #1

A producer sells a product at the point where its Marginal Cost (MC) is equal to or less than the Marginal Revenue (MR) or th
Answer #2
(60-55)+(45-55)=-10+5=- $5
Answer #3
http://www.transtutors.com/questions/tts-shoe-company-a-domestic-shoe-company-distributes-running-shoes-and-tennis-shoes--130326.htm
Answer #4

Answer

Domestic sell

Production cost

Profit

Remarks

Both shoes (95+95)=190

60+45=105

190-105=$85

He would sell to domestic retailer

Sell to Chinese buyer

55+55=110

60+45=105

110-105=$5

He would not sell to Chinese buyer

a.     Should the shoe company sell running shoes to the Chinese retailer?

Answer No shoe company will not sell to Chinese retailer

b.    Why or why not?

Answer : profit is less comparison to domestic selling. Or Chinese retailer is purchasing company good lower price than in the domestic market and it is less than the marginal cost of production so shoe company will not sell below the marginal cost condition. Loss will incur to company if he only sell running shoe (95-55)=$45

c.     Should the chose company sell tennis shoes to the Chinese retailer?

Answer No shoe company will not sell to Chinese retailer

d.    Why or why not?

Answer : profit is less comparison to domestic selling. Or Chinese retailer is purchasing company good lower price than in the domestic market and no doubt it is paying more than the marginal cost of production even though shoe company will not sell because loss will incur to company if he only sell tennis shoes (95-55)=$45

        e.             Answer

Selling price

Total cost

Profit

Remarks

95*45000=4275000

95*45000=4275000

                   8550000

45000*60=2700000

45000*45=2025000

                    4725000

8550000-4725000=3825000

(15000+15000)*55=

1650000

15000*60=900000

15000*45=675000

                    1575000

1650000-1575000=75000

               

        

        Shoe company will not sell to Chinese retailer.